Physics 335—Thermal Physics Homework Assignment #5 Due Wednesday, October 9, 2024, 11:40xam

Ch 4:

Problems 4.3, 4.10, and 4.14 (20 pts each).

Hints: For problem 4.10, make some reasonable estimates for the temperature of your kitchen and the temperature inside the refrigerator. (Manufacturers typically recommend about 0^{∘}F for the freezer portion of a household refrigerator, and about 39^{∘}F for the refrigerator portion. The freezer temperature is the relevant one for determining the coefficient of performance.)